# Diffusion-empowered AutoPrompt MedSAM

### v1

@@ 4.3 Ablation Study
4.3.1 Effects of Components in AutoMedSAM
We conducted systematic ablation experiments on the Chest Xray Masks and Labels dataset to evaluate the specific contribution of each component to the method’s performance. As shown in Table 4, removing different components led to varying degrees of performance degradation. When both diffusion processing and uncertainty-aware joint optimization were excluded from the training process, the DSC and NSD of the model dropped to 88.62% and 89.19%, respectively, approximately a 7% decrease compared to the complete model, indicating the significant importance of these two components. Further analysis revealed that the absence of either component alone also caused substantial performance degradation. Additionally, removing the MSE loss and the pre-trained knowledge transferred from MedSAM (used to assist in training the diffusion-based class prompt encoder) had the second most significant impact on performance after the exclusion of diffusion processing or joint optimization. Finally, the absence of other components also led to varying degrees of performance reduction. In conclusion, optimal segmentation performance can only be achieved by incorporating all components during training.
Table 4: Ablation study on different components during AutoMedSAM training. The gray row represents our method.
Defusion Processing Joint Optimization CE Dice SD MSE DSC(%$\uparrow$ NSD(%$\uparrow$
@@ 4.3.1 Effects of Components in AutoMedSAM
4.3.2 Limitations of Manual Prompts
@@
As discussed in Sec. 1, manually defined prompt boxes significantly constrain the segmentation accuracy of the model. To further investigate this phenomenon, we examined the impact of prompt boxes with varying offsets on the segmentation results. The experimental results, shown in Table 5, indicate that segmentation accuracy improves as the boundaries of the prompt box approach the target object. However, the accuracy is not the best when the box aligns perfectly with the target boundary. Fig. 6 illustrates the segmentation masks under different prompt box configurations. The figure reveals that overly large prompt boxes include multiple segmentable objects, leading to misidentification by the SAM-core model, as it struggles to determine the specific organ to segment. Conversely, overly small prompt boxes drive the model to search for deeper internal differences within the box, which reduces segmentation accuracy. To address these challenges, we proposed the class prompt method, which incorporates semantic information into the segmentation process. This approach eliminates errors caused by unstable manual prompts, simplifies the segmentation procedure, and enhances the model’s robustness.
@@ 5 Conclusion
In this paper, we proposed AutoMedSAM, a fully automated SAM-based medical image segmentation solution that demonstrates high effectiveness. AutoMedSAM consists of three components: an image encoder, a diffusion-based class prompt encoder, and a mask decoder. The diffusion-based class prompt encoder extracts relevant information from the image embedding based on the prompted class, and through two separate decoder branches, generates sparse and dense prompt embeddings. This design eliminates the need for manual prompts and additional prompt generators, achieving an end-to-end workflow that enhances system robustness and broadens application scenarios and user groups. Additionally, we introduced an uncertainty-aware joint optimization process, enabling the model to adaptively determine the weights of different loss functions to improve generalization capability. During joint optimization, we supervise both intermediate processes and final outputs to transfer the extensive knowledge of MedSAM to AutoMedSAM effectively. These strategies not only improve training efficiency but also significantly enhance segmentation accuracy. However, our current method has only been trained on the datasets mentioned in this paper, and it does not yet fully meet the precision segmentation requirements for all organs and lesions. In the future, we will explore more lightweight solutions and further optimize the model on larger-scale medical datasets to enhance its practicality and applicability.

### v2

@@ 4.4 Ablation Study
4.4.1 Effects of Prompt Branch
The diffusion-based prompt encoder structure shows that we designed two branches to enhance its performance. To determine their impact on the final segmentation accuracy, we conducted experiments on each branch separately (shown in Table 5). During the experiments, one branch was deactivated when the other was activated. The missing prompt embeddings are initialized in accordance with the approach adopted by MedSAM. Table 5 shows that AutoMedSAM achieves optimal performance when both branches are activated simultaneously. When the Sparse branch is activated, the NSD decreases by 4.616%, whereas activating the Dense branch results in a 1.09% decrease. This demonstrates that the Dense branch plays a complementary role in enhancing boundary details. And the embeddings from both branches mutually reinforce each other, ultimately contributing to optimal prompt.
Table 5: Ablation study on the prompt branch using the AbdomenCT-1K dataset.
Prompt Branch DSC(%$\uparrow$ NSD(%$\uparrow$
Dense Branch 93.752 94.058
Sparse Branch 93.256 90.532
Dense + Sparse Branch(Ours) 94.580 95.148
4.4.2 Effects of Diffusion Processing
Following the characteristics of diffusion models, AutoMedSAM incorporates class prompts conditionally during the forward diffusion process, enabling the model to learn to emphasize target-specific features across varying noise levels, thereby enhancing its representation of complex anatomical structures. We performed ablation experiments to verify the diffusion process’s effect on the prompts. Table 6 shows the experimental results. From the table, it can be noticed that when the diffusion process is disabled, the DSC and NSD are reduced by 5.379% and 5.173% respectively. This proves that diffusion processing is the key to AutoMedSAM’s robust generalization in complex clinical settings.
Table 6: Ablation study on the diffusion processing using the Chest-XML dataset.
Diffusion Processing DSC(%$\uparrow$ NSD(%$\uparrow$
91.562 92.194
96.941 97.367
4.4.3 Effects of Uncertainty-aware Joint Optimization
Table 7: Ablation study on the Uncertainty-aware joint optimization using the Chest-XML dataset.
Joint Optimization CE Dice SD MSE DSC(%$\uparrow$ NSD(%$\uparrow$
@@ 4.4.3 Effects of Uncertainty-aware Joint Optimization
To evaluate the contribution of the proposed uncertainty-aware joint optimization with each loss, we conducted a series of ablation experiments. As shown in Table 7, we progressively disabled the different loss components (e.g., CE, DC, SD, and MSE), and observed the resulting changes in segmentation performance. The full strategy with all four losses included in the uncertainty-aware weighting framework achieved the highest performance with DSC and NSD metrics of 96.941% and 97.367%, respectively. When the uncertainty weighting was removed and fixed weights were used, the DSC and NSD metrics decreased by 3.582% and 3.254%, respectively, indicating that the adaptive weighting mechanism is very effective in reconciling conflicting loss contributions. Among the individual loss components, removing the MSE loss (used to transfer MedSAM pre-training knowledge) resulted in the most significant decrease (from 96.941% to 92.573%), indicating its important role in knowledge transfer. Disabling CE, DC, or SD loss also led to significant degradation, highlighting the complementary nature of pixel-based, region-based, and distribution-based targets when capturing fine anatomical structures. Overall, uncertainty-aware joint optimization can effectively balance multiple loss functions, and improve segmentation accuracy and generalization.
4.5 Limitations of Manual Prompts
As discussed in Sec. 1, manually sketching an explicit prompt significantly constrains the segmentation accuracy of the model. To further investigate this phenomenon, we explored the impact of prompt boxes with varying offsets on the segmentation results. The experimental results indicate that segmentation accuracy improves as the boundaries of the prompt box approach the target object(shown in Table 8). However, the accuracy is not the best when the box aligns perfectly with the target boundary. Fig. 6 illustrates the segmentation masks under different prompt box configurations. The figure reveals that overly large prompt boxes include multiple segmentable objects, leading to misidentification by the SAM-core model, as it struggles to determine the specific organ to segment. Conversely, overly small prompt boxes drive the model to search for deeper internal differences within the box, which reduces segmentation accuracy. To address these challenges, we proposed the class prompt method, which incorporates semantic information into the segmentation process. This approach eliminates errors caused by unstable manual prompts, simplifies the segmentation procedure, and enhances the model’s robustness.
@@ 5 Conclusion
In this paper, we present AutoMedSAM, an end-to-end interactive model for medical image segmentation, designed to address the limitations of manual prompting and the lack of semantic annotation in SAM. By introducing a diffusion-based class prompt encoder, AutoMedSAM eliminates the need for labor-intensive prompts (e.g., points, boxes, or scribbles), and instead utilizes class indices to generate prompt embeddings. The diffusion process conditionally integrates class prompts throughout the generation of embeddings, progressively enhancing the model’s ability to capture the structural and semantic characteristics of organs or lesions, thereby improving the controllability and quality of the generated prompts. The dual-branch architecture further refines both local and global features, boosting segmentation detail and stability. Additionally, we propose an uncertainty-aware joint optimization strategy that adaptively integrates pixel-based, region-based, and distribution-based losses, improving generalization and robustness across diverse medical modalities. Extensive experiments demonstrate that AutoMedSAM achieves state-of-the-art performance, and also shows strong cross-dataset generalization. With the help of AutoMedSAM, doctors and researchers can quickly and accurately analyze medical images in time-sensitive clinical settings, enabling more timely diagnosis and treatment. In the future, we will optimize it on larger-scale medical datasets to enhance its practicality and clinical applicability.